OpenStudy (anonymous):

Given a sequence of Bernoulli trials with n=7 and p=0.8 , find the binomial probabilities for the k values given below. k=0 k=1 k=2 k=3

OpenStudy (aum):

Probability of "k" successes = \[\large \left(\begin{matrix}n \\ k\end{matrix}\right)p^k(1-p)^{n-k} \]Plug in the values.

OpenStudy (aum):

\[ \large \left(\begin{matrix}n \\ k\end{matrix}\right) = \frac{n!}{(n-k)! * k!} \]

OpenStudy (anonymous):

okay let me see